# A quasi-optimal upper bound for induced paths in sparse graphs

### v1

@@ 3.2 Index-trees and index-barriers
Suppose that $i\in{\sf T}(b^{-})$. By construction every $i$ in $B(\ell)$ in fact appears in an occurrence of $B(b^{-})$, and every such occurrence of $B(b^{-})$ does not contain the letter $b$ and is preceded and followed by $b$ in $B(\ell)$. Hence we have $W=b\circ B(b^{-})\circ b$.
Otherwise let $k$ be the ancestor of $b$ (with possibly $k=b$) such that $i\in{\sf T}(k^{+})$. Again, note that by construction each such $i$ in $B(\ell)$ lies in an occurrence of $B(k)$. We argue that $W$ must be factor of $B(k)$. Indeed, note that if $W$ is not a factor of $B(k)$ then by Observation 3.5 it contains the parent of $k$. But then, by construction, as $i$ comes from ${\sf T}(k^{+})$, $W$ must contain $B(k^{-})$ as well. However $b\in B(k^{-})$, a contradiction. So $W$ is a factor of
$B(k)=k\circ B(k^{-})\circ k\circ B(k^{+})\circ k\circ B(k^{-})\circ k,$
with every occurrence of $i$ lying in $B(k^{+})$, and occurrences of $b$ lying in $k\circ B(k^{-})\circ k$. We conclude that $W$ starts with the last occurrence of $b$ in $k\circ B(k^{-})\circ k$ and ends with the first occurrence of $b$ in $k\circ B(k^{-})\circ k$.
@@
By Proposition 3.7, since $i<b$, $i$ either lies in ${\sf T}(k_{1})$ for $k_{1}=b^{-}$, or in ${\sf T}(k_{1}^{+})$ for some ancestor $k_{1}$ of $b$ with possibly $k_{1}=b$. Since $j<b$, we obtain analogously that $j$ lies in ${\sf T}(k_{2})$ for $k_{2}=b^{-}$, or in ${\sf T}(k_{2}^{+})$ for an ancestor $k_{2}$ of $b$.
Let us argue that $i$ and $j$ lie in the same subtree, i.e., that $k_{1}=k_{2}$. Suppose that it is not the case, and consider a minimum factor containing $i$ and $j$ but avoiding $b$ in $B(\ell)$. Note that such a factor exists since $i\operatorname{{\sf reach}}_{b}j$. By Observation 3.5, this factor contains the indices of the shortest path connecting $k_{1}$ and $k_{2}$ in ${\sf T}$. This rules out the cases $k_{1}=b$ and $k_{2}=b$ and implies that the factor contains at least two ancestors of $b$. Let $k$ be the lowest such ancestor in ${\sf T}$. Note that since one of $i,j$ lies in ${\sf T}(k^{+})$, we have $b\in{\sf T}(k^{-})$. Then, by definition of $B(\ell)$, the factor must contain all the indices of ${\sf T}(k^{-})$, including $b$, a contradiction. Hence $k_{1}=k_{2}$ and the statement follows. ∎
@@
$\operatorname{{\sf reach}}^{+}_{b}(i)$ =$=\{j:j\operatorname{{\sf reach}}_{b}i\ \text{and}\ j>i\},$
$\operatorname{{\sf reach}}^{-}_{b}(i)$ =$=\{j:j\operatorname{{\sf reach}}_{b}i\ \text{and}\ j<i\}.$
@@ 3.3 Ribbed-tree and node-barriers
Let $s,t$ be two adjacent nodes of ${\sf S}_{\ell}$, with $s\prec t$, and let $i,j$ be their respective ranks. Then either $i\geq 2$, in which case we have $j=i-1$, or $i=1$, in which case $j$ may be any index in $\{1,\dots,\ell-1\}$; see Section 3.1.
Figure 4: The construction of the barriers in the two cases where $4=i\geq 2$ (left) and $i=1$ (right), for $\ell=7$ and with ${\sf T}$ being a complete binary tree as in Figure 3. The ranks of the nodes are indicated by colored integers. The shaded areas correspond to the zones, with the zone of 666 partially represented on the left for better readability: it should be understood that it contains the two other represented zones as well. The square node (right) represents the extra subdivision $z$ in the case where $i=1$. Note that the obtained subdivided nodes are considered in the zone of their parent node $s$.
Let us first describe the modifications to perform when $i\geq 2$. These modifications are better understood accompanied with Figure 4. Let $p=|B(i,j)|-2$. If $p=0$, then $i$ and $j$ are adjacent in ${\sf T}$. In this case we do not modify the edge $st$. Otherwise, we subdivide it $p$ times, where each obtained node will represent an internal index of $B(i,j)$. Formally, let $y_{1},\dots,y_{p}$ be the obtained subdivided vertices, in this order from $s$ to $t$. Let $c_{1},\dots,c_{p}\in\left\{1,\dots,\ell\right\}$ be the indices such that $B(i,j)=ic_{1}\cdots c_{p}j$. For every $1\leq k\leq p$ we add the edge $y_{k}s_{k}$ where $s_{k}={\sf zone}\text{-}{\sf ancestor}_{s}(c_{k})$, i.e., we connect $y_{k}$ to the node of ${\sf S}_{\ell}$ of rank $c_{k}$ whose zone contains $s$. Note that distinct such $y_{k}$ may be adjacent to a same node $s_{k}$, if they represent a same index of the index-barrier. We point that these modifications are performed indifferently if $t$ is a left child of $s$, or if it is a right child.
If $i=1$ we proceed analogously as in the first case, except that we consider the index-barrier from $j$ to $i$ instead, since $j\geq i$, and subdivide the edge once more with the final node being adjacent to its ancestor of rank $j$. More formally, we proceed as follows. If $i=j=1$, let $p=0$. Otherwise let $p=|B(i,j)|-2$ and note that $p\geq 0$. We subdivide $p+1$ times the edge $st$. Let $z,y_{1},\dots,y_{p}$ be the obtained vertices, in order from $t$ to $s$ this time, i.e., from bottom to top in the skeleton-tree. If $p\geq 1$, let $c_{1},\dots,c_{p}\in\left\{1,\dots,\ell\right\}$ be the indices such that $B(j,1)=jc_{1}\cdots c_{p}1$. For every $1\leq k\leq p$ we add the edge $y_{k}s_{k}$ where $s_{k}={\sf zone}\text{-}{\sf ancestor}_{s}(c_{k})$. Then we add the edge $zs^{\prime}$ for $s^{\prime}={\sf zone}\text{-}{\sf ancestor}_{s}(j)$.
@@ 3.4 The blow-up of a ribbed-tree
In the previous section we described the concept of a ribbed-tree. The purpose of the edges added in the construction of the ribbed-tree is to ensure that there is no long induced path, because our final goal is to construct a graph where all the induced paths are short. However, we also want this graph to have a long path and to be 2-degenerate. Therefore in this section we describe an operation to transform a ribbed-tree into a graph that contains a Hamiltonian path and is 222-degenerate (the proof that it is indeed the case is given in Section 3.5). This operation, that we call the blow-upmargin: blow-up , is defined as follows. It is similar to the “blow-up operation” described in [DR24] but differs on the way ribs are handled. These modifications are better understood accompanied with Figures 5 and 6.
@@
  • Recall that for every blocking-node $x$ of the barrier a rib $e$ was connecting it to an ancestor node $s^{\prime}$. To represent $e$, we create three ribs $e_{1},e_{2},e_{3}$ with $x_{1}$ connecting $e_{1}$ to the bottom triangle-vertex of the triangle $K^{s^{\prime}}$, $e_{2}$ connecting $x_{2}$ to the right-endpoint of the top-edge of this triangle, and $e_{3}$ connecting $x_{3}$ to the left-endpoint. We call left barrier-pathmargin: left barrier-path the obtained path, and left barriermargin: left barrier the left barrier-path together with its incident ribs.
@@
Figure 6: A partial representation of the graph $G$ where ${\sf T}$ is a complete binary tree on three indices. The rank of triangle-vertices are depicted within the triangles. For better readability, not all ribs are depicted; it should be understood that the root triangle sends ribs to any blocking-vertex of rank 333, in purple in the picture.
In the following, to avoid confusion between $G$ and the other auxiliary structures, we will exclusively refer to the elements of $V(G)$ as vertices. Note that for every triangle-vertex $v$ in $G$, there is a node $t$ of ${\sf S}_{\ell}$ such that $v$ belongs to the triangle $K^{t}$. We denote this node by $\pi(v)$margin: $\pi(v)$ Similarly, if $x$ is a representative then $\pi(x)$ denote the blocking node of ${\sf R}$ it originates from. We say that $K^{t}$ is a root trianglemargin: root triangle if $t$ a root.
@@
We conclude the section with the next definition and observation on the sequence of ranks in barrier-paths. Recall that the collapse of a word is the contraction of repeated consecutive entries; see Section 2. The last remark of the below observation follows from Observation 3.12 and the presence of hopping ribs.
@@
Moreover, if $Q^{\prime}$ is the subpath of $Q$ consisting of blocking-vertices, then its trace does not contain $i$ and $i-1$ if $i\geq 2$, and it always contains $j$ if $i=1$.
3.5 Hamiltonian path and degeneracy
@@
Consequently, an analysis identical to the one conducted in [DR24, Lemmas 4.1 and 4.3] yields the two following lemmas, whose proofs are omitted.
@@
For any index-tree ${\sf T}$ the graph $G({\sf T})$ is 222-degenerate.
@@ 4 Properties of induced paths
  • ${P}{v}$ is the prefix of $P$ ending in $v$;
  • ${v}{P}$ is the suffix of $P$ starting in $v$;
  • ${u}{P}{v}$ is the subpath from $u$ to $v$; and
@@
If $u,v$ are two consecutive vertices in ${\sf triangle}\text{-}{\sf seq}(P)$, then the trace of ${u}{P}{v}$ is a factor of $B(\ell)$.
@@
If $u^{\prime}$ lies in a barrier, then either $P$ crosses the full barrier, in which case it reaches $v$ in the attained triangle, or $P$ contains a rib attached to the barrier, in which case $v$ is the top endpoint of that rib. In the first case, by Observation 3.19, the trace of ${u}{P}{v}$ is $B(i,j)$ with $i$ and $j$ being the respective ranks of $u$ and $v$, or of $v$ and $u$, depending on whether ${\sf rank}(u)>{\sf rank}(v)$ or ${\sf rank}(u)<{\sf rank}(v)$. Hence this trace is a factor of $B(\ell)$. In the other case, note that ${u}{P}{v}$ has visited a part of $B(i,j)$, starting from ${\sf rank}(u)$ to an occurrence of ${\sf rank}(v)$. Hence the trace of ${u}{P}{v}$ is a factor of $B(\ell)$ as desired.
In the remaining case where $u^{\prime}$ is the endpoint of a rib followed by $P$, it lies in a barrier separating two cliques $K^{s}$ and $K^{t}$, whose trace is $B(i,j)$ for some $i,j\in\{1,\dots,\ell\}$. Then $P$ may either reach one of these two triangles, or reach another index in the barrier and take a corresponding rib. The same arguments as above shows that the trace of ${u}{P}{v}$ is a factor of $B(\ell)$ in these cases. ∎
Notice that once an induced path that visits a triangle vertex $u$, and after if it visits vertices of other ranks, then it may not contain the bottom endpoint of a rib incident to $u$. Additionally, we will argue that if an induced path has intersected a rib on its bottom endpoint, then it will not be able to contain other ribs of that rank within the same zone. We formalize the conditions that should be verified by $P$ in order to verify these properties.
@@
A triangle-vertex $v$ of $P$ burnsmargin: to burn an index $i\in\left\{1,\dots,\ell\right\}$ if $i$ is a ghost rank, or if ${P}{v}$ contains the bottom endpoint of a rib of index $i$ incident to some $u\in{\sf zone}\text{-}{\sf ancestor}_{v}(i)$.
@@
Figure 7: Three possible situations for a vertex $u$ to be correct when its rank is 444 (left) and 333 (right); in this example it is assumed that ${\sf T}$ is a complete binary tree as in Figure 3. In red and blue are represented different possibilities for ${P}{u}$. In blue, the index-lock of $u$ is 777: note that ${P}{u}$ intersects every rib of rank in $\operatorname{{\sf reach}}^{+}_{7}(4)=\{5,6\}$ (left), or every rib of rank in $\operatorname{{\sf reach}}^{+}_{7}(3)=\emptyset$ (right). In red, the index-lock of $u$ is 666, and 777 is a ghost rank: it is the only index in $\operatorname{{\sf reach}}^{+}_{6}(3)$.
The key step of our proof is showing that every triangle-vertex of $P$ is correct, which we use to characterize the way $P$ behave between consecutive vertices in ${\sf triangle}\text{-}{\sf seq}(P)$. For this, we will rely on the following property which expresses a notion of “confinement” for vertices following a correct triangle-vertex.
@@
  • ${u}{P}$ does not contain a vertex of rank $b$ in ${\sf zone}_{u}(b)$; and
  • ${u}{P}$ does not contain a triangle-vertex of rank $i\in\operatorname{{\sf reach}}^{+}_{b}(a)$ in ${\sf zone}_{u}(b)$, and in particular not a rib of that rank contained in ${\sf zone}_{u}(b)$.

### v2

@@ 3.2 Index-trees and index-barriers
Suppose that $i\in{\sf T}(b^{-})$. By construction every $i$ in $B(\ell)$ in fact appears in a factor of the form $B(b^{-})$, and every such occurrence of $B(b^{-})$ does not contain the letter $b$ and is preceded and followed by $b$ in $B(\ell)$. Hence we have $W=b\circ B(b^{-})\circ b$.
Otherwise let $k$ be the ancestor of $b$ (with possibly $k=b$) such that $i\in{\sf T}(k^{+})$. Again, note that by construction each occurrence of $i$ in $B(\ell)$ lies in a factor of the form $B(k)$. We argue that $W$ must be factor of $B(k)$. Indeed, note that if $W$ is not a factor of $B(k)$ then by Observation 3.5 it contains the parent of $k$.
But then, by construction, $W$ contains either a suffix or a prefix of
$B(k)=k\circ B(k^{-})\circ k\circ B(k^{+})\circ k\circ B(k^{-})\circ k.$
As $i$ appears only in $B(k^{+})$ and $b$ appears in $k\circ B(k^{-})$, we get that $b$ is an internal letter of $W$, a contradiction.
So $W$ is indeed a factor of $B(k)$. From the definition of $B(k)$ as recalled just above, we conclude that $W$ starts with the last occurrence of $b$ in the prefix $k\circ B(k^{-})\circ k$ and ends with the first occurrence of $b$ in the suffix $k\circ B(k^{-})\circ k$.
@@
By Proposition 3.7, since $i<b$, $i$ either lies in ${\sf T}(b^{-})$, or in ${\sf T}(k_{1}^{+})$ for some ancestor $k_{1}$ of $b$ with possibly $k_{1}=b$. Since $j<b$, we obtain analogously that $j$ lies in ${\sf T}(b^{-})$, or in ${\sf T}(k_{2}^{+})$ for an ancestor $k_{2}$ of $b$.
If $i,j\in{\sf T}(b^{-})$ we are done. Below we consider the other cases. Consider a minimum factor $W$ of $B(\ell)$ containing $i$ and $j$ but avoiding $b$, which exists since $i\operatorname{{\sf reach}}_{b}j$. By Observation 3.5, this factor contains the indices of the shortest path connecting $k_{1}$ and $k_{2}$ in $\sf T$}.
In the case where $i\in{\sf T}(b^{-})$ and $j\in{\sf T}(k_{2}^{+})$ for some ancestor $k_{2}$ of $b$ (and the symmetric case where the roles of $i$ and $j$ are reversed), $W$ contains $b$ since it lies on the $i$-$j$ path of $\sf T$}, a contradiction.
In the case where $i\in V({\sf T}(k_{1}^{+}))$ and $j\in V({\sf T}(k_{2}^{+}))$ for some ancestors $k_{1}^{+},k_{2}^{+}$ of $b$, we may assume by symmetry that $k_{1}^{+}$ is lower than $k_{2}^{+}$. Then $b\in V({\sf T}(k_{1}^{-})$. If $k_{1}\neq k_{2}$, by definition of $B(\ell)$ the factor $W$ contains all the indices of ${\sf T}(k_{1}^{-})$, including $b$, a contradiction. Hence $k_{1}=k_{2}$ and the statement follows. ∎
@@
$\operatorname{{\sf reach}}^{+}_{b}(i)$ $=\{j:j\operatorname{{\sf reach}}_{b}i\ \text{and}\ j>i\},$
$\operatorname{{\sf reach}}^{-}_{b}(i)$ $=\{j:j\operatorname{{\sf reach}}_{b}i\ \text{and}\ j<i\}.$
@@ 3.3 Ribbed-tree and node-barriers
Let $s,t$ be two adjacent nodes of $\sf S$$_{\ell}$, with $s\prec t$, and let $i,j$ be their respective ranks. Then either $i\geq 2$, in which case we have j=i1j=i-1, or $i=1$, in which case $j$ may be any index in $\{1,\dots,\ell-1\}$; see Remark 3.1.
Figure 4: The construction of the barriers in the two cases where $4=i\geq 2$ (left) and $i=1$ (right), for $\ell=7$ and with $\sf T$} being a complete binary tree as in Figure 3. The ranks of the nodes are indicated by colored integers. The shaded areas correspond to the zones, with the zone of 66 partially represented on the left for better readability: it should be understood that it contains the two other represented zones as well. The square node (right) represents the extra subdivision $z$ in the case where $i=1$. Note that the obtained subdivided nodes are considered in the zone of their parent node $s$.
Let us first describe the modifications to perform when $i\geq 2$. These modifications are better understood accompanied with Figure 4. Let p=|B(i,j)|2p=|B(i,j)|-2. If $p=0$, then $i$ and $j$ are adjacent in $\sf T$}. In this case we do not modify the edge $st$. Otherwise, we subdivide it $p$ times, where each obtained node will represent an internal index of $B(i,j)$. Formally, let $y_{1},\dots,y_{p}$ be the obtained subdivided vertices, in this order from $s$ to $t$. Let $c_{1},\dots,c_{p}\in\left\{1,\dots,\ell\right\}$ be the indices such that $B(i,j)=ic_{1}\cdots c_{p}j$. For every $1\leq k\leq p$ we add the edge $y_{k}s_{k}$ where $s_{k}={\sf zone}\text{-}{\sf ancestor}_{s}(c_{k})$, i.e., we connect $y_{k}$ to the node of $\sf S$$_{\ell}$ of rank $c_{k}$ whose zone contains $s$. Note that distinct such $y_{k}$ may be adjacent to a same node $s_{k}$, if they represent a same index of the index-barrier. We point that these modifications are performed indifferently if $t$ is a left child of $s$, or if it is a right child.
If $i=1$ we proceed analogously as in the first case, except that we consider the index-barrier from $j$ to $i$ instead, since $j\geq i$, and subdivide the edge once more with the final node being adjacent to its ancestor of rank $j$. More formally, we proceed as follows. If $i=j=1$, let $p=0$. Otherwise let p=|B(i,j)|2p=|B(i,j)|-2 and note that $p\geq 0$. We subdivide $p+1$ times the edge $st$. Let $z,y_{1},\dots,y_{p}$ be the obtained vertices, in order from $t$ to $s$ this time, i.e., from bottom to top in the skeleton-tree. If $p\geq 1$, let $c_{1},\dots,c_{p}\in\left\{1,\dots,\ell\right\}$ be the indices such that $B(j,1)=jc_{1}\cdots c_{p}1$. For every $1\leq k\leq p$ we add the edge $y_{k}s_{k}$ where $s_{k}={\sf zone}\text{-}{\sf ancestor}_{s}(c_{k})$. Then we add the edge $zs^{\prime}$ for $s^{\prime}={\sf zone}\text{-}{\sf ancestor}_{s}(j)$.
@@ 3.4 The blow-up of a ribbed-tree
In the previous section we described the concept of a ribbed-tree. The purpose of the edges added in the construction of the ribbed-tree is to ensure that there is no long induced path, because our final goal is to construct a graph where all the induced paths are short. However, we also want this graph to have a long path and to be 2-degenerate. Therefore in this section we describe an operation to transform a ribbed-tree into a graph that contains a Hamiltonian path and is $2$-degenerate (the proof that it is indeed the case is given in Section 3.5). This operation, that we call the blow-upmargin: blow-up , is defined as follows. It is similar to the “blow-up operation” described in [DR24] but differs on the way ribs are handled. These modifications are better understood accompanied with Figures 5 and 6.
@@
  • Recall that for every blocking-node $x$ of the barrier, a rib $e$ was connecting it to an ancestor node $s^{\prime}$. To represent $e$, we create three ribs $e_{1},e_{2},e_{3}$ with $x_{1}$ connecting $e_{1}$ to the bottom triangle-vertex of the triangle $K^{s^{\prime}}$, $e_{2}$ connecting $x_{2}$ to the right-endpoint of the top-edge of this triangle, and $e_{3}$ connecting $x_{3}$ to the left-endpoint. We call left barrier-pathmargin: left barrier-path the obtained path, and left barriermargin: left barrier the left barrier-path together with its incident ribs.
@@
Figure 6: A partial representation of the graph $G$ where $\sf T$} is a complete binary tree on three indices. The rank of triangle-vertices are depicted within the triangles. For better readability, not all ribs are depicted; it should be understood that the triangle at the root sends ribs to any blocking-vertex of rank $3$, in purple in the picture.
In the following, to avoid confusion between $G$ and the other auxiliary structures, we will exclusively refer to the elements of $V(G)$ as vertices. Note that for every triangle-vertex $v$ in $G$, there is a node $t$ of $\sf S$$_{\ell}$ such that $v$ belongs to the triangle $K^{t}$. We denote this node by $\pi(v)$margin: $\pi(v)$ . Similarly, if $x$ is a representative then $\pi(x)$ denote the blocking node of $\sf R$} it originates from.
@@
We conclude the section with the next definition and observation on the sequence of ranks in barrier-paths. Recall that the collapse of a word is the contraction of repeated consecutive entries; see Section 2. The last remark of the following observation is a consequence of Observation 3.12 and the presence of hopping ribs.
@@
Moreover, if $Q^{\prime}$ is the subpath of $Q$ consisting of blocking-vertices, then its trace does not contain $i$ and i1i-1 if $i\geq 2$, and it always contains $j$ if $i=1$.
3.5 Traceability and degeneracy
@@
Consequently, an analysis similar to the one conducted in [DR24, Lemmas 4.1 and 4.3] yields the two following lemmas, whose proofs are included for the sake of completeness.
@@
Proof.
Let $\ell=V({\sf T})$ (i.e. $\sf T$} is a $\ell$-index tree) and $G=G({\sf T})$. For ever node $s$ of $\sf S$$_{\ell}$ we denote by $G_{\downarrow s}$ the subgraph of $G$ induced by all the triangle vertices of some $K^{t}$ where $t$ is a descendant of $s$ in $\sf S$$_{\ell}$ (including the case $t=s$) as well as all barrier-paths between these triangles.
We will prove by induction on the structure of the graph the following statement: for every node $s$ of $\sf S$$_{\ell}$, the graph $G_{\downarrow s}$ has a Hamiltonian path from the right to the left top triangle vertices of $K^{s}$. This implies the statement of the lemma when $s$ is the root of $S_{\ell}$.
The base case is when $s$ is a leaf of $\sf S$$_{\ell}$. In this case $G_{\downarrow s}$ is a triangle and the desired property clearly holds.
Suppose now that $s$ is not a leaf and that the property holds both for it left child $t_{L}$ and its right child $t_{R}$. Then a Hamiltonian path of $G_{\downarrow s}$ can be obtained as follows: starting from the right top triangle vertex of $K^{s}$, we follow the barrier-path leading to the right top triangle vertex of $K^{t_{R}}$, then we follow the Hamiltonian path of $G_{\downarrow t_{R}}$ provided by the induction hypothesis and end up at the left top triangle vertex of $K^{t_{R}}$, then we follow the barrier-path leading to the bottom triangle vertex of $K^{s}$ followed by that leading to the right top triangle vertex of $K^{t_{L}}$, we follow the Hamiltonian path of $G_{\downarrow t_{L}}$ given by the induction hypothesis and conclude with the barrier-path ending at the left top triangle vertex of $K^{s}$. ∎
@@
For any index-tree $\sf T$} the graph $G({\sf T})$ is $2$-degenerate.
Proof.
[…]
  1. 1.
    ($u\in K^{s}$ or $u$ is an internal vertex of some barrier-path above $K_{s}$) and ($v\in K^{t}$ or $v$ is in an internal vertex of some barrier-path above $K^{t}$), or
  2. 2.
    $u$ and $v$ belong to the same barrier-path between $K^{s}$ and $K^{t}$ (in whih case $t$ is a child node of $s$ in $\sf T$$_{\ell}$) and $K^{s}$ is closer from $u$ than from $v$; or
  3. 3.
    ($u,v\in K^{s}$ or $u,v\in K^{t}$) and $v$ is the bottom triangle vertex.
[…]
  • If $u$ is the interior vertex of a barrier path above $K^{t}$ for some $t\in{\sf S}_{\ell}$, it has 3 neighbors in $G$, by construction (two on the barrier-path and one via a rib). However the neighbor $v$ of $u$ on the barrier-path that is the closest to $K^{t}$ is such that $u\sqsubseteq v$ according to item 2 above, hence $v\notin V(H)$ (by maximality of $u$) and $u$ indeed has at most two neighbors.
  • If $u$ is a triangle vertex in some $K^{s}$ then in $G$ it is has neighbors that are endpoints of ribs, however by construction every such vertex belongs to a barrier paths above $K^{t}$ for some $t\in{\sf S}_{\ell}$ with $s\prec t$, so as above such neighbor do not appear in $H$. The same is true of neighbors in the barrier paths above $K^{t}$ when $t$ is a child of $s$. The possible remaining neighbors are the two other vertices of $K^{s}$ if $u$ is the bottom triangle vertex, or, if $u$ is a top triangle vertex, the other top triangle vertex and a vertex in the barrier path above. In both cases $u$ has degree at most 2 in $G$.
Recall that triangle vertices and interior vertices of barrier-path partition the vertex set of $G$, so we considered all cases. Hence $G$ is 2-degenerate. ∎
@@ 4 Properties of induced paths
  • Pv{P}{v} is the prefix of $P$ ending in $v$;
  • vP{v}{P} is the suffix of $P$ starting in $v$;
  • uPv{u}{P}{v} is the subpath from $u$ to $v$; and
@@
If $u,v$ are two consecutive vertices in ${\sf triangle}\text{-}{\sf seq}(P)$, then the trace of uPv{u}{P}{v} is a factor of $B(\ell)$.
@@
If $e$ lies in a barrier, then either $P$ crosses the full barrier, in which case it reaches $v$ in the attained triangle, or $P$ contains a rib attached to the barrier, in which case $v$ is the top endpoint of that rib. In the first case, by Observation 3.19, the trace of uPv{u}{P}{v} is $B(i,j)$ with $i$ and $j$ being the respective ranks of $u$ and $v$, or of $v$ and $u$, depending on whether ${\sf rank}(u)>{\sf rank}(v)$ or ${\sf rank}(u)<{\sf rank}(v)$. Hence this trace is a factor of $B(\ell)$. In the other case, note that uPv{u}{P}{v} has visited a part of $B(i,j)$, starting from ${\sf rank}(u)$ to an occurrence of ${\sf rank}(v)$. Hence the trace of uPv{u}{P}{v} is a factor of $B(\ell)$ as desired.
In the remaining case where $e$ is a rib, $u^{\prime}$ lies in a barrier separating two triangle $K^{s}$ and $K^{t}$, whose trace is $B(i,j)$ for some $i,j\in\{1,\dots,\ell\}$. Then $P$ may either reach one of these two triangles, or reach another index in the barrier and take a corresponding rib. The same arguments as above shows that the trace of uPv{u}{P}{v} is a factor of $B(\ell)$ in these cases. ∎
Notice that if an induced path visits a triangle vertex $u$ without following the ribs attached to it, then it will not contain the bottom endpoints of duch ribs. Additionally, we will argue that if an induced path has intersected a rib on its bottom endpoint, then it will not be able to contain other ribs of that rank within the same zone. We formalize the conditions that should be verified by $P$ in order to verify these properties.
@@
A triangle-vertex $v$ of $P$ burnsmargin: to burn an index $i\in\left\{1,\dots,\ell\right\}$ if $i$ is a ghost rank, or if Pv{P}{v} contains the bottom endpoint of a rib of index $i$ incident to some $u\in{\sf zone}\text{-}{\sf ancestor}_{v}(i)$.
@@
Figure 7: Three possible situations for a vertex $u$ to be correct when its rank is 44 (left) and 33 (right); in this example it is assumed that $\sf T$} is a complete binary tree as in Figure 3. In red and blue are represented different possibilities for Pu{P}{u}. In blue, the index-lock of $u$ is 77: note that Pu{P}{u} intersects every rib of rank in $\operatorname{{\sf reach}}^{+}_{7}(4)=\{5,6\}$ (left), or every rib of rank in $\operatorname{{\sf reach}}^{+}_{7}(3)=\emptyset$ (right). In red, the index-lock of $u$ is $6$, and $7$ is a ghost rank: it is the only index in $\operatorname{{\sf reach}}^{+}_{6}(3)$.
The key step of our proof is showing that every triangle-vertex of $P$ is correct, which we use to characterize the way $P$ behaves between consecutive vertices in ${\sf triangle}\text{-}{\sf seq}(P)$. For this, we will rely on the following property which expresses a notion of “confinement” for vertices following a correct triangle-vertex.
@@
  • uP{u}{P} does not contain a vertex of rank $b$ in ${\sf zone}_{u}(b)$; and
  • uP{u}{P} does not contain a triangle-vertex of rank $i\in\operatorname{{\sf reach}}^{+}_{b}(a)$ in ${\sf zone}_{u}(b)$, and in particular not a rib of that rank contained in ${\sf zone}_{u}(b)$.